Question: A triangle has a 12 degree angle and a 73 degree angle, find the measure of the third angle
I get 85? Not sure though, appreciate any help in advance.

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

12 + 73 = 85 for the 2 angles
The total of the angles of all triangles is 180 degs
180 - 85 = 95 degs, the 3rd angle.

Answer by rfer(16322)

Your answer should be 95.
180-12-73=95